OpenStudy (theedwardsfamily):

HELP ME PLZ!!!! Medal +Fan What is the missing exponent? (106) ____= 10 -18

Nnesha (nnesha):

the question is \[\huge\rm (10^6)^{??}=10^{-18}\] use `^` for exponent

OpenStudy (bossvideogamer21):

i honestly don't know

OpenStudy (anonymous):

oh that formatting helps alot! haha

Nnesha (nnesha):

now remember the exponent rule \[\huge\rm (x^m)^n=x^{m ·n}\] you should multiply the missing exponent with 6 to get -18 exponent

Nnesha (nnesha):

what would you multiply by 6 to get -18 ?

OpenStudy (theedwardsfamily):

ok 6*(-3) = -18

Nnesha (nnesha):

looks good

Nnesha (nnesha):

or there is another way to find missing exponent let missing exponent = x \[\huge\rm (10^6)^{x}=10^{-18}\] \[\huge\rm 10^{6x}=10^{-18}\]bases are the same so you cancel them out and then solve for x

OpenStudy (theedwardsfamily):

yea so -3 would be the answer

OpenStudy (theedwardsfamily):

I get it now

Nnesha (nnesha):

good

OpenStudy (theedwardsfamily):

Thanks @Nnesha !